OpenStudy (anonymous):

A floor plan is given for the first floor of a new house. One inch represents 6 feet. If a small closet is 1/4 in by 1/3 in on the floor plan, what are the actual dimensions? A.1.36 ft by 1.5 ft B. 1.75 ft by 2 ft C. 1.5 ft by 2 ft D. 2 ft by 2 ft

OpenStudy (wolf1728):

An inch on the floor plan is 6 feet in real life. So 1/4 inch = 6/4 feet = 1.5 feet 1/3 inch = 6/3 feet = 2 feet So the answer is C